40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4614     Entry time: Tue May 3 15:48:26 2011
Author: Larisa Thorne 
Type: Update 
Category: Electronics 
Subject: laser temperature control LPF, final version! 

This is a continuation of this

 The low pass filter is finally acceptable, and its Bode graph is below (on a ~3Hz frequency span that shows the cutoff frequency is at 0.1Hz)

Attachment 1: LPF100k.2.jpg  19 kB  | Hide | Hide all
LPF100k.2.jpg
ELOG V3.1.3-